A person of average density of 980 kg/m3 floats in seawater of density 1025 kg/m3. What can you determine using this information? Determine everything.
Answer #1

D = 1025kg/m^2

d = 980kg/m^2

Let height of person submerged in water is h

And H be the total height of person.

Hence h/H = 980/1025 = 0.956

=) h = (0.956)H

That means 95.6% of person height is inside seawater.

And 100-95.6 = 4.4% is out from seawater because seawater is more denser than person .